# Context-free graphs and their transition groups

### v2

@@ 2.1. Basic facts about morphisms of inverse graphs
Let $\varphi_{1},\varphi_{2}$ be two morphisms from the inverse graph $\Gamma^{(1)}$ to the inverse graph $\Gamma^{(2)}$, then $\varphi_{1}=\varphi_{2}$ if and only if $\varphi_{1}(x)=\varphi_{2}(x)$ for some $x\in V(\Gamma)$.
@@
Any walk $x\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}y$ is mapped by $\varphi_{1},\varphi_{2}$ into the two walks
$\varphi_{1}(x)\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}\varphi_{1}(y)\mbox{ and }\varphi_{2}(x)\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}\varphi_{2}(y)$
and so by the determinism of $\Gamma^{(2)}$ and condition $\varphi_{1}(x)=\varphi_{2}(x)$ we deduce $\varphi_{1}(y)=\varphi_{2}(y)$. Since $\Gamma^{(1)}$ is connected, we have that $\varphi_{1},\varphi_{2}$ agree on $V(\Gamma^{(1)})$, hence by the determinism of $\Gamma^{(2)}$, they also agree on $E(\Gamma^{(1)})$. ∎
Another interesting feature of inverse graphs, and particularly rooted inverse graphs, is the relationship between accepted languages and morphisms. This connection is contained in the following proposition belonging to folklore, see for instance [24, Proposition 1].
@@
Let $(\Gamma,x_{0}),(\Lambda,y_{0})$ be two rooted inverse graphs, then $L(\Gamma,x_{0})\subseteq L(\Lambda,y_{0})$ if and only if there is a morphism $\psi:\Gamma\rightarrow\Lambda$ with $\psi(x_{0})=y_{0}$. Moreover, $L(\Gamma,x_{0})=L(\Lambda,y_{0})$ if and only if there is an isomorphism $\psi:\Gamma\rightarrow\Lambda$.
Note that by the previous proposition, we have that, up to isomorphism, there is a unique rooted inverse graph accepting the language $L(\Gamma,x_{0})$. The set of automorphisms of the inverse graph $\Gamma$ will be denoted by $\operatorname{Aut}(\Gamma)$. This is a group acting on $\Gamma$ in a canonical way. In case the quotient graph $\Gamma\setminus\operatorname{Aut}(\Gamma)$ is finite, we say that $\Gamma$ is quasi-transitive. Observe that this last condition is equivalent to the finiteness of $V(\Gamma)\setminus\operatorname{Aut}(\Gamma)$. Any morphism $\varphi:\Gamma\to\Lambda$ between two inverse graphs induces a bijections $\varphi:\operatorname{Star}(\Gamma,v)\to\operatorname{Star}(\Gamma^{\prime},\varphi(v))$ between the star sets in the obvious ways. Thus, morphisms of inverse graphs are immersions of graphs as defined by Stallings [25]. In particular, if the inverse graphs are complete, by Lemma 2.3 every morphism between two complete inverse graphs is a covering, that is, an immersion of graphs $f:\Gamma\to\Lambda$ so that $f$ maps the sets $V(\Gamma)$ and $E(\Gamma)$ respectively onto $V(\Lambda)$ and $E(\Lambda)$. One central feature of coverings is the possibility of lifting walks, this is contained in the following characterization.
@@
Let $\varphi:\Gamma^{\prime}\to\Gamma$ be a morphism of inverse graphs. Then, for each vertex $v\in V(\Gamma)$ and vertex $v^{\prime}\in\varphi^{-1}(v)$, every walk $p$ in $\Gamma$ with $\iota(p)=v$ lifts to a unique walk $p^{\prime}$ starting at $v^{\prime}$, i.e., $\varphi(p^{\prime})=p$.
3. Context-free grammars, pushdown automata, and configurations graphs
There are several ways to define a context-free language, either via the notion of grammar, or using the notion of pushdown automaton. For the sake of completeness, we briefly recall the main notions that are used in the paper, and we refer the reader to [18] for further details.
@@ 3. Context-free grammars, pushdown automata, and configurations graphs
A context-free grammar is a tuple $\mathcal{G}=(V,\Sigma,P,S)$ where $V$ is a finite set of non-terminal symbols (or variables), $\Sigma$ is the finite alphabet of terminal symbols that is disjoint from $V$, $S\in V$ is the starting symbol, and $P\subseteq V\times(V\cup\Sigma)^{*}$ is a finite set called productions.
Usually production are denoted by $X\Rightarrow_{\mathcal{G}}\alpha$, $\alpha\in(V\cup\Sigma)^{*},X\in V$. We extend $\Rightarrow_{\mathcal{G}}$ by putting $\alpha X\gamma\Rightarrow_{\mathcal{G}}\alpha\beta\gamma$ whenever $X\Rightarrow_{\mathcal{G}}\beta\in P$, and we denote by $\Rightarrow_{\mathcal{G}}^{*}$ the reflexive and transitive closure of $\Rightarrow_{\mathcal{G}}$. The language generated by the grammas $\mathcal{G}$ is $L(S)=\{u\in\Sigma^{*}:S\Rightarrow^{*}_{\mathcal{G}}u\}$, and in this case we say that it is a context-free language.
[…]
In what follows, we will concentrate our attention on deterministic pushdown automata that are real-time, so without $1$-moves. Pushdown automata without $1$-moves have configurations graphs without edges labeled with the empty word $1$, so it is the natural framework when considering configurations graphs that are inverse $\Sigma$-digraphs. It is clear that to have $\Gamma(M)$ that is a deterministic $\Sigma$-digraph, we need to assume $M$ deterministic. The natural question is when $\Gamma(M)$ is also co-deterministic, since in this case we may transform $\Gamma(M)$ into an inverse graph $\Gamma(M)^{\pm}$ by simply adding for every edge $(q,\xi x)\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}(p,\xi\gamma)$ the reverse one $(p,\xi\gamma)\smash{\stackrel{{\scriptstyle\text{\tiny{$a^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}(q,\xi x)$ (see Remark 2.2). In particular, it is of interest to see when $\Gamma(M)^{\pm}$ is the configurations graph of some deterministic PDA. We shall see that this happens if $M$ has a reverse transition function, so $M$ is a reversible pushdown automaton. A reversible pushdown automaton $M=(Q,\Sigma,\Theta,\delta,q_{0},\perp,F)$ is characterized by having a reverse transition function $\delta_{R}:Q\times\Sigma\times\Theta\to Q\times\Theta^{*}$ such that is able to locally reverse the computation, i.e.,
[…]
We remark that the initial assumption that $M$ is real-time is not strict from the language point of view since by [19, Theorem 5] for every reversible pushdown automaton $N$ there is a real-time reversible PDA $M$ such that $L(M)=L(N)$. To deal with inverse graphs, a new class of PDAs that is more suitable to consider languages on involutive closed alphabet $\Sigma\cup\Sigma^{-1}$ in [24] has been introduced.
Definition 3.2 (Inverse PDA and inverse-context-free languages).
A real-time deterministic PDA $M=(Q,\Sigma\cup\Sigma^{-1},\Theta,\delta,q_{0},\perp,F)$ is called inverse if $\delta$ may be partitioned into two functions: $\delta^{+}$ which is the restriction of $\delta$ to $Q\times\Sigma\times\Theta$ and $\delta^{-}$ which is the restriction of $\delta$ to $Q\times\Sigma^{-1}\times\Theta$ such that $\delta^{+}$ is the reverse transition of $\delta^{-}$ and vice versa in the following sense:
[…]
A language $L\subseteq(\Sigma\cup\Sigma^{-1})^{*}$ is called inverse-context-free if $L=L(M)$ for some inverse PDA $M$ on $\Sigma\cup\Sigma^{-1}$.
It is clear from the definition and the fact that $\delta^{+},\delta^{-}$ are functions, that the configurations graph $\Gamma(M)$ is inverse if and only if $M$ is an inverse PDA. For an invertible PDA $M$ let us set $M^{+}=(Q,\Sigma,\Theta,\delta^{+},q_{0},\perp,F)$. It is straightforward to see that a PDA $M$ is inverse if and only if $M^{+}$ is reversible.
3.1. Inverse context-free graphs
The machine point of view that we have seen before can be reformulated into a more graph-theoretic setting via the notion of context-free graph. The starting point is the notion of context-free group given in [22]. In their paper, Muller and Schupp called a finitely generated group $G$ presented by $\langle A\mid\mathcal{R}\rangle$ context-free if the word problem $WP(G;A)$ is a context-free language. With our notation, this language coincides with the set $L(\operatorname{Cay}(G;A),\mathds{1}_{G})$ of words labeling a circuit starting and ending at the group identity.
@@ 3.1. Inverse context-free graphs
A rooted inverse graph $(\Gamma,x_{0})$ is called context-free if $L(\Gamma,x_{0})$ is a context-free language.
This definition is equivalent to a more geometric one that involves the notion of end-cones. We briefly recall it. Henceforth, we denote by $d(x,y)$ the usual distance in the inverse graph $\Gamma$, and put $\|v\|_{x_{0}}=d(x_{0},v)$. In [23] Muller and Schupp introduced the central notion of end-cone together with the idea of end-isomorphism of such subgraphs. For a vertex $v$ with $\|v\|_{x_{0}}=n$, the end-cone at $v$ is the (inverse) subgraph $\Gamma(v,x_{0})$ that is the connected component of $\Gamma\setminus D_{n-1}(x_{0})$ containing $v$, and we denote by $\Delta(v,x_{0})=\{y\in\Gamma(v,x_{0}):\|y\|_{x_{0}}=n\}$ the set of frontier vertices (or vertices) of $\Gamma(v,x_{0})$. An end-isomorphism $\psi:\Gamma(v,x_{0})\to\Gamma(w,x_{0})$ between the end-cones $\Gamma(v,x_{0}),\Gamma(w,x_{0})$ is an isomorphism of inverse graphs with the property of preserving their frontier vertices: $\psi(\Delta(v,x_{0}))=\Delta(w,x_{0})$. We sometimes write $\Gamma(v,x_{0})\simeq_{end}\Gamma(w,x_{0})$ to emphasize the existence of an end-isomorphism between the two end-cones.
@@
The rooted graph $(\Gamma,x_{0})$ is called a context-free graph in the sense of Muller-Schupp if up to end-isomorphism there are finitely many end-cones, i.e., there are inverse graphs $\Gamma_{1},\ldots,\Gamma_{N}$ such that for any end-cone $\Gamma(v,x_{0})$ of $\Gamma$ there is an end-isomorphism $\psi:\Gamma(v,x_{0})\to\Gamma_{j}$ for some $j\in[1,N]$.
In [23] it is shown that the graph $(\Gamma(M),x_{0})$ of the instantaneous descriptions of a pushdown automaton is a context-free graph, and conversely given a context-free graph $(\Gamma,x_{0})$ there is a pushdown automaton $M$ such that $(\Gamma(M),x_{0})$ is isomorphic to $(\Gamma,x_{0})$. The following characterization relates all the notions that we have seen so far.
@@
For an inverse graph $\Gamma$ the following conditions are equivalent:
@@
    for any non-empty finite set $F\subseteq V(\Gamma)$ the language $L(\Gamma,x_{0},F)$ is context-free;
@@
    the language $L(\Gamma,x_{0})$ is context-free, hence $(\Gamma,x_{0})$ is context-free;
@@
    there exists an inverse PDA $M$ with initial state $q_{0}$ and initial stack symbol $\perp$ such that the rooted configurations graph $(\Gamma(M),(q_{0},\perp))$ is isomorphic to $(\Gamma,x_{0})$;
  5. (5)
    for any vertex $y_{0}\in V(\Gamma)$, the rooted graph $(\Gamma,y_{0})$ is context-free in the sense of Muller-Schupp.
4. Transition groups of inverse graphs and co-context-free groups
In this section, we introduce a new subclass of co-context-free groups. For an $A$-digraph $\Gamma$, the set of everywhere-circuits is
(1) $\mathcal{L}(\Gamma)=\{u\in A^{*}:\forall p\in V(\Gamma)\,\mbox{ there is a circuit }p\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p\mbox{ in }\Gamma\}=\bigcap_{p\in V(\Gamma)}L(\Gamma,p)$
It is easy to see that if $\Gamma$ is an inverse $A$-digraph that is also complete, then $\mathcal{L}(\Gamma)$ is a normal subgroup of $F_{A}$. In this setting, we give the following definition.
@@ 4. Transition groups of inverse graphs and co-context-free groups
For a complete inverse $A$-digraph $\Gamma$, each letter $a\in A$ defines a permutation on the set of vertices $a:V(\Gamma)\to V(\Gamma)$. The transition group $\mathcal{G}(\Gamma)$ of the inverse graph $\Gamma$ is the group generated by composing all these permutations.
It is easy to see that $\mathcal{G}(\Gamma)$ may be equivalently obtained by considering the quotient of the free group $F_{A}$ generated by $A$ by the normal subgroup $\mathcal{L}(\Gamma)$. There is a natural right action of $\mathcal{G}(\Gamma)$ on the set of vertices $V(\Gamma)$: for any $g\in\mathcal{G}(\Gamma)$ and vertex $v\in V(\Gamma)$, $v\cdot g\in V(\Gamma)$ is the vertex for which there is a walk $v\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}(v\cdot g)$ in $\Gamma$ for any $w\in A^{*}$ that represents $g$. We sometimes also put $v\cdot w=v\cdot g$ for any $w\in A^{*}$ representing $g$ in $\mathcal{G}(\Gamma)$.
@@
By taking $\Gamma$ either as a finite $n$-cycle or an infinite line and by suitably labeling all edges either by $a$ or $a^{-1}$ from a single-letter alphabet $A=\{a\}$, we can turn $\Gamma$ into a complete inverse graph so that its transition group is readily seen to be either isomorphic to $\mathbb{Z}/n\mathbb{Z}$ or $\mathbb{Z}$, so that all these groups are transition groups.
@@
The transition group is the syntactic monoid of the language $L=L(\Gamma,x_{0})$ (where $x_{0}$ is any chosen vertex of $\Gamma$). Indeed, it is not difficult to see that if one considers the syntactic congruence $\sim_{\Gamma}$ on $A^{*}$ defined by $u\sim_{L}v$ if for all $x,y\in A^{*}$ $xuy\in L$ if and only if $xvy\in L$, then since $\Gamma$ is inverse, we have that $A^{*}/_{\sim_{L}}$ is isomorphic to $\mathcal{G}(\Gamma)$. If we remove the completeness condition, the syntactic monoid is an inverse monoid. For a general introduction on syntactic monoids of a formal language see [13].
Note that the rooted graph $(\Gamma,p)$ may be naturally identified with the Schreier graph of the stabilizer subgroup $H=\mathcal{G}(\Gamma)_{p}=\{g\in\mathcal{G}(\Gamma):g(p)=p\}$. This fact follows from Proposition 2.5 and the equality $L(\Gamma,p)=L(\operatorname{Sch}(H,A),H)$.
We actually want to focus on a slightly more general framework: we consider a finite collection of inverse complete $A$-digraphs $\{\Gamma^{(1)},\ldots,\Gamma^{(k)}\}$ and we see that, as before, we can consider the permutations induced by the letters of $A$. In this way, we speak about the transition group
$\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)}):=F_{A}/_{\mathcal{L}(\Gamma^{(1)},\ldots,\Gamma^{(k)})}$
with $\mathcal{L}(\Gamma^{(1)},\ldots,\Gamma^{(k)})=\bigcap_{i=1}^{k}\mathcal{L}(\Gamma^{(i)})$.
It will be useful to see a transition group over a finite union of graphs as a the subgroup of a direct product of transition groups over a single graph. Namely, we have the following lemma.
@@
If $\Gamma^{(1)},\ldots,\Gamma^{(k)}$ are inverse $A$-graphs that are complete, then the transition group $\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ is isomorphic to a subgroup of $\mathcal{G}(\Gamma^{(1)})\times\ldots\times\mathcal{G}(\Gamma^{(k)})$.

### v3

@@ 2.1. Basic facts about morphisms of graphs
Let $\varphi_{1},\varphi_{2}$ be two morphisms from the graph $\Gamma^{(1)}$ to the graph $\Gamma^{(2)}$, then $\varphi_{1}=\varphi_{2}$ if and only if $\varphi_{1}(x)=\varphi_{2}(x)$ for some $x\in V(\Gamma)$.
@@
For any walk $x\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}y$, the images under $\varphi_{1}$ and $\varphi_{2}$ are $\varphi_{1}(x)\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}\varphi_{1}(y),\varphi_{2}(x)\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}\varphi_{2}(y)$. Determinism of $\Gamma^{(2)}$ and $\varphi_{1}(x)=\varphi_{2}(x)$ imply $\varphi_{1}(y)=\varphi_{2}(y)$. Since $\Gamma^{(1)}$ is connected, $\varphi_{1}$ and $\varphi_{2}$ coincide on all vertices, and hence also on all edges. ∎
Another interesting feature of graphs, and particularly rooted graphs, is the relationship between accepted languages and morphisms. This connection is contained in the following proposition belonging to folklore, see for instance [27, Proposition 1].
@@
Let $(\Gamma,x_{0}),(\Lambda,y_{0})$ be two rooted graphs, then $L(\Gamma,x_{0})\subseteq L(\Lambda,y_{0})$ if and only if there is a morphism $\psi:\Gamma\rightarrow\Lambda$ with $\psi(x_{0})=y_{0}$. Moreover, $L(\Gamma,x_{0})=L(\Lambda,y_{0})$ if and only if there is an isomorphism $\psi:\Gamma\rightarrow\Lambda$.
Note that by the previous proposition, we have that, up to isomorphism, there is a unique rooted graph accepting the language $L(\Gamma,x_{0})$. The set of automorphisms of the graph $\Gamma$ will be denoted by $\operatorname{Aut}(\Gamma)$. This is a group acting on $\Gamma$ in a canonical way. In case the quotient graph $\Gamma\setminus\operatorname{Aut}(\Gamma)$ is finite, we say that $\Gamma$ is quasi-transitive. Observe that this last condition is equivalent to the finiteness of $V(\Gamma)\setminus\operatorname{Aut}(\Gamma)$. Any morphism $\varphi:\Gamma\to\Lambda$ between two graphs induces a bijections $\varphi:\operatorname{Star}(\Gamma,v)\to\operatorname{Star}(\Gamma^{\prime},\varphi(v))$ between the star sets in the obvious ways. Thus, morphisms of graphs are immersions of graphs as defined by Stallings [29]. In particular, if the graphs are complete, by Lemma 2.3 every morphism between two complete $A$-graphs is a covering, that is, an immersion of graphs $f:\Gamma\to\Lambda$ so that $f$ maps the sets $V(\Gamma)$ and $E(\Gamma)$ respectively onto $V(\Lambda)$ and $E(\Lambda)$. One central feature of coverings is the possibility of lifting walks, this is contained in the following characterization.
@@
Let $\varphi:\Gamma^{\prime}\to\Gamma$ be a covering of graphs. Then, for each vertex $v\in V(\Gamma)$ and vertex $v^{\prime}\in\varphi^{-1}(v)$, every walk $p$ in $\Gamma$ with $\iota(p)=v$ lifts to a unique walk $p^{\prime}$ starting at $v^{\prime}$, i.e., $\varphi(p^{\prime})=p$.
3. Context-free languages and graphs
There are several ways to define a context-free language, either via the notion of grammar, or using the notion of pushdown automaton. For the sake of conciseness, we briefly recall the main notions that are used in the paper, and we refer the reader to [19] for further details.
@@ 3. Context-free languages and graphs
A context-free grammar is a tuple $(V,\Sigma,P,S)$ where $V$ is a finite set of non-terminal symbols (or variables), $\Sigma$ is the finite alphabet of terminal symbols that is disjoint from $V$, $S\in V$ is the starting symbol, and $P\subseteq V\times(V\cup\Sigma)^{*}$ is a finite set called productions.
Usually production are denoted by $X\Rightarrow\alpha$, $\alpha\in(V\cup\Sigma)^{*},X\in V$. We extend $\Rightarrow$ by putting $\alpha X\gamma\Rightarrow\alpha\beta\gamma$ whenever $X\Rightarrow\beta\in P$, and we denote by $\Rightarrow^{*}$ the reflexive and transitive closure of $\Rightarrow$. The language generated by the grammar $(V,\Sigma,P,S)$ is $L(S)=\{u\in\Sigma^{*}:S\Rightarrow^{*}u\}$, and in this case we say that it is a context-free language. A $L$ language with a context-free complement $\Sigma^{*}\setminus L$ is called co-context-free, we will denote the class of context-free languages as CoCF.
In their paper, Muller and Schupp called a finitely generated group $G$ presented by $\langle A\mid\mathcal{R}\rangle$ context-free if the word problem $WP(G;A)=\{\,w\in A^{*}\mid\pi(w)=\mathds{1}_{G}\}$ is a context-free language, where $\pi:A^{*}\to G$ denotes the natural projection. With our notation, this language coincides with the set $L(\operatorname{Cay}(G;A),\mathds{1}_{G})$ of words labeling a circuit starting and ending at the group identity.
@@
A rooted graph $(\Gamma,x_{0})$ is called context-free if $L(\Gamma,x_{0})$ is a context-free language.
This definition is equivalent to a more geometric one that involves the notion of end-cones. We briefly recall it. Henceforth, we denote by $d(x,y)$ the usual distance in the graph $\Gamma$, and put $\|v\|_{x_{0}}=d(x_{0},v)$. In [26] Muller and Schupp introduced the central notion of end-cone together with the idea of end-isomorphism of such subgraphs. For a vertex $v$ with $\|v\|_{x_{0}}=n$, the end-cone at $v$ is the subgraph $\Gamma(v,x_{0})$ that is the connected component of $\Gamma\setminus D_{n-1}(x_{0})$ containing $v$, and we denote by $\Delta(v,x_{0})=\{y\in\Gamma(v,x_{0}):\|y\|_{x_{0}}=n\}$ the set of frontier vertices of $\Gamma(v,x_{0})$, that is, the vertices of $\Gamma(v,x_{0})$ at distance $n$ from the root $x_{0}$. An end-isomorphism $\psi:\Gamma(v,x_{0})\to\Gamma(w,x_{0})$ between the end-cones $\Gamma(v,x_{0}),\Gamma(w,x_{0})$ is an isomorphism of graphs with the property of preserving their frontier vertices: $\psi(\Delta(v,x_{0}))=\Delta(w,x_{0})$.
@@
The rooted graph $(\Gamma,x_{0})$ is called a context-free graph in the sense of Muller-Schupp if up to end-isomorphism there are finitely many end-cones, i.e., there are graphs $\Gamma_{1},\ldots,\Gamma_{N}$ such that for any end-cone $\Gamma(v,x_{0})$ of $\Gamma$ there is an end-isomorphism $\psi:\Gamma(v,x_{0})\to\Gamma_{j}$ for some $j\in[1,N]$.
The previous two notions are equivalent:
@@
For a graph $\Gamma$ the following conditions are equivalent:
@@
    for any non-empty finite set $F\subseteq V(\Gamma)$ the language $L(\Gamma,x_{0},F)$ is deterministic context-free;
@@
    the language $L(\Gamma,x_{0})$ is deterministic context-free, hence $(\Gamma,x_{0})$ is context-free;
@@
    for any vertex $y_{0}\in V(\Gamma)$, the graph $(\Gamma,y_{0})$ is context-free in the sense of Muller-Schupp.
The languages accepted by context-free graphs can be characterized more precisely through the construction of suitable pushdown automata, called inverse PDA. In [27], such languages are called inverse-context-free, and we will use this term throughout the paper.
4. Transition groups of graphs and co-context-free groups
In this section, we introduce a new subclass of co-context-free groups. For an $A$-graph $\Gamma$, the set of everywhere-circuits is
(1) $\mathcal{L}(\Gamma)=\{u\in A^{*}:\forall p\in V(\Gamma)\,\mbox{ there is a circuit }p\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}p\mbox{ in }\Gamma\}=\bigcap_{p\in V(\Gamma)}L(\Gamma,p)$
It is easy to see that if $\Gamma$ is an $A$-graph that is also complete, then $\mathcal{L}(\Gamma)$ is a normal subgroup of $F_{A}$. The following is a key notion in our paper.
@@ 4. Transition groups of graphs and co-context-free groups
For a complete $A$-graph $\Gamma$, each letter $a\in A$ defines a permutation on the set of vertices $a:V(\Gamma)\to V(\Gamma)$. The transition group $\mathcal{G}(\Gamma)$ of the graph $\Gamma$ is the group generated by composing all these permutations.
Remark 4.2.
Note that this definition is a specialization of the well-known notion of transition monoid for an automaton to the case of inverse automata. This is also isomorphic to the syntactic monoid of the language $L=L(\Gamma,x_{0})$ (where $x_{0}$ is any chosen vertex of $\Gamma$). Indeed, it is not difficult to see that if one considers the syntactic congruence $\sim_{\Gamma}$ on $A^{*}$ defined by $u\sim_{L}v$ if for all $x,y\in A^{*}$ $xuy\in L$ if and only if $xvy\in L$, then since $\Gamma$ is inverse, we have that $A^{*}/_{\sim_{L}}$ is isomorphic to $\mathcal{G}(\Gamma)$. If we remove the completeness condition, the syntactic monoid is an inverse monoid, see [14, 28].
The group $\mathcal{G}(\Gamma)$ may be equivalently obtained by considering the quotient of the free group $F_{A}$ generated by $A$ by the normal subgroup $\mathcal{L}(\Gamma)$. There is a natural right action of $\mathcal{G}(\Gamma)$ on the set of vertices $V(\Gamma)$: for any $g\in\mathcal{G}(\Gamma)$ and vertex $v\in V(\Gamma)$, $v\cdot g\in V(\Gamma)$ is the vertex for which there is a walk $v\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$w$}}}{{\relbar\joinrel\longrightarrow}}}(v\cdot g)$ in $\Gamma$ for any $w\in A^{*}$ that represents $g$. We sometimes also put $v\cdot w=v\cdot g$ for any $w\in A^{*}$ representing $g$ in $\mathcal{G}(\Gamma)$.
@@
For any group $G$ on the symmetric generating set $A$, it is straightforward to check that the transition group of the Cayley graph $\operatorname{Cay}(G;A)$ is isomorphic to $G$. In particular, if $\Gamma$ is either a finite $n$-cycle or an infinite line and by suitably labeling all edges either by $a$ or $a^{-1}$ from a single-letter alphabet $A=\{a\}$, we can turn $\Gamma$ into a complete graph so that its transition group is readily seen to be either isomorphic to $\mathbb{Z}/n\mathbb{Z}$ or $\mathbb{Z}$, so that all these groups are transition groups.
More in general, the rooted graph $(\Gamma,p)$ may be naturally identified with the Schreier graph of the stabilizer subgroup $H=\mathcal{G}(\Gamma)_{p}=\{g\in\mathcal{G}(\Gamma):g(p)=p\}$. This fact follows from Proposition 2.5 and the equality $L(\Gamma,p)=L(\operatorname{Sch}(H,A),H)$. On the other hand, given a group $G$ generated by $A$ and a subgroup $H$, the Schreier graph of $H$, i.e. the graph which vertices are right coset of $H$ and direct edges are labeled by elements in $A$, is an example of a complete $A$-graph. In view of this, one can provide an algebraic characterization of transition groups of a graph.
Proposition 4.4.
A group $G$ is a transition group for a complete $A$-graph $\Gamma$ if and only $G$ admits a core-free subgroup $H$.
Proof.
Let $p$ be any vertex of $\Gamma$ and set $H:=G_{p}$, the stabilizer of $p$. Since for any $g\in G$ the conjugate $G_{p}^{g}=G_{p\cdot g}$, the normal core of $H$ (the intersection of all conjugates of $H$) coincides with the intersection of all vertex stabilizers in $\Gamma$. By the definition of a transition group, this intersection is trivial. Conversely, consider the Schreier graph $\operatorname{Sch}(H,A)$ of $H$. As previously discussed, it is a complete $A$-digraph on which $G$ acts naturally. Let $w$ be a word in $\mathcal{L}(\operatorname{Sch}(H,A),H)$. By definition, $w$ represents an element of $G$ that stabilizes all vertices of the Schreier graph; that is, $Hg\cdot\overline{w}=Hg$ for all $g\in G$. Therefore, $\overline{w}$ belongs to the normal core of $H$, which is trivial, so $w\in WP(G,A)$. This shows that $\mathcal{L}(\operatorname{Sch}(H,A),H)\subseteq WP(G,A)$. The reverse inclusion is straightforward, hence $G$ is indeed the transition group of $\operatorname{Sch}(H,A)$. ∎
@@
By the previous proposition we immediately deduce that the growth of the group $\mathcal{G}(\Gamma)$ is bounded below by the growth of the graph $\Gamma$.
We actually aim to work in a slightly more general setting. Let $\{\Gamma^{(1)},\ldots,\Gamma^{(k)}\}$ be a finite collection of complete $A$-graphs. As before, we can consider the permutations of vertices induced by the letters of $A$. This allows us to define the transition group associated with the collection as $\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)}):=F_{A}/_{\mathcal{L}(\Gamma^{(1)},\ldots,\Gamma^{(k)})}$, where
$\mathcal{L}(\Gamma^{(1)},\ldots,\Gamma^{(k)}):=\bigcap_{i=1}^{k}\mathcal{L}(\Gamma^{(i)})$
denotes the set of words acting trivially on all the graphs in the collection. It is convenient to view a transition group over a finite union of graphs as a subgroup of the direct product of the corresponding transition groups over single graphs. More precisely, we have the following fact.
@@
Let $\Gamma^{(1)},\ldots,\Gamma^{(k)}$ be complete $A$-graphs. Then $\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})\hookrightarrow\mathcal{G}(\Gamma^{(1)})\times\ldots\times\mathcal{G}(\Gamma^{(k)})$, i.e., the transition group of the union is isomorphic to a subgroup of the direct product of the individual transition groups.